OpenStudy (anonymous):

Find all the values of c that make x^2+cx+64 a perfect square.

OpenStudy (apoorvk):

so if we check for the form a^2 + 2ab + b^2, a^ = x^2 b^2 = 64, so b= +8, or -8 so, cx = 2.8.x or 2.(-8).x so, c = 16 or -16

jhonyy9 (jhonyy9):

16 will be right sure

OpenStudy (apoorvk):

-16 as well

jhonyy9 (jhonyy9):

yes -/+ 16 are right too